# Partial Resolution of the Erdős-Straus, Sierpiński, and Generalized Erdős-Straus Conjectures Using New Analytical Formulas

### v1

@@ Abstract
This article proposes a unified analytical approach leading to a partial resolution of the Erdős–Straus, Sierpiński conjectures, and their generalization. We introduce two new analytical formulas under specific conditions of divisibility and the existence of perfect squares. Under these conditions, the formulas verify the conjectures even for very large numerical values. Moreover, our method reduces the problem to the search for a suitable perfect square, thereby opening the way to a complete proof of these conjectures. Notably, our second formula significantly improves upon Mordell’s work by demonstrating analytically the conjecture in the majority of cases where Mordell’s approach fails. Furthermore, these formulas are highly versatile, as they provide, under the established conditions, a systematic method to decompose any fraction $\frac{a}{n}$ into a sum of three Egyptian fractions. In conclusion, we present open questions and conjectures to the mathematical community regarding the generalization of these formulas.
@@ 1 Introduction
The Erdős-Straus conjecture, formulated in 1948 by Paul Erdős and Ernst G. Strauss, states that for every natural number $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
@@
Louis Mordell demonstrated that if $n$ is not congruent to $1,11^{2},13^{2},17^{2},19^{2},23^{2}$ modulo 840, then the conjecture holds. Furthermore, numerical verifications have confirmed its validity up to $n=10^{17}$.
@@
The Polish mathematician Wacław Sierpiński also investigated this conjecture and posed a similar question: Is a similar decomposition possible if 4 is replaced by 5? He conjectured that for every natural number $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
@@
Finally, following the formulation of the Erdős-Straus conjecture, several mathematicians explored the possibility of generalizing this Egyptian fraction representation. This led to the generalized Erdős-Straus conjecture, which states that for every $a\geq 4$ and every $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
@@
We will show that our two analytical formulas also verify this conjecture.
@@ 2 The Erdős-Straus Conjecture
One of the methods that is known nowadays is that if $n$ is even integer , meaning $n=2k$ with $k\in\mathbb{N}^{*}$. In this case, it is well established that:
@@
In my new approach, I will demonstrate how this result can be naturally recovered when $n$ is even. Before introducing the general method that allowed me to establish two formulas verifying the conjecture under a certain condition, I will first illustrate my process by analytically testing the conjecture for some values of $n$.
My method is based on the structural analysis of equation (1). It is a simple yet powerful and general approach, applicable to even numbers, odd numbers, and prime numbers alike.
@@ 2.1 Case n = 2 (Even Number)
I choose not to simplify $\frac{4}{2}$ immediately in order to maintain the general structure of the method. However, one can simplify this fraction directly, and by following the same steps that I will present, the same result is always obtained.
@@
Since we are searching for solutions in $\mathbb{N}^{*}$, it is necessary that $24x-$ 2 be strictly positive, that is, $04x-2>$ 0, which gives $x\geq\lfloor\frac{1}{2}\rfloor+1=1$.
For the rest of the document, the notation $\lfloor x\rfloor$ will always denote the integer part of $x$ to avoid any confusion. We then obtain the equation:
$yz(4x-2)=2x(y+z).$
@@ 2.2 Parameterization Technique
Let $a,b,c,d\in\mathbb{R}$ such that:
@@
This parametric representation holds for any $t$ and allows rewriting $(dt)b=(bt)d$, which verifies the initial equation. This technique will be crucial for my results.
@@ 2.3 Application to Our Equation
This parameterization is valid for all $t\in\mathbb{N}^{*}$, as we are looking for solutions in $\mathbb{N}^{*}$.
@@
For $y$ and $z$ to be integers, $\Delta$ must be a perfect square, and $24x-$ 2 must have the same parity as $\sqrt{\Delta}$.
@@ 2.4 Analysis of the Method
One of the key advantages of our analytical approach is its applicability not only to even numbers but also to odd numbers and prime numbers. Unlike classical methods-which have relied primarily on brute-force techniques that merely test the existence of solutions without offering insight into the problem’s structure—our method derives solutions directly from the inherent logical structure of the underlying Diophantine equation. This provides a general, rigorous framework for addressing the Erdős-Straus, Sierpiński, and generalized Erdős-Straus conjectures. Moreover, the elegance and clarity of our approach make it a valuable pedagogical tool for teaching methods of solving Diophantine equations, as it not only yields the solutions but also deepens our understanding of the arithmetic properties involved
@@ 2.5 Case $n=3$ (odd and prime Number)
For $t=1$, we find a perfect square $12111^{2}=$ 121, yielding $y=1$ and $z=12$.
@@ 4 Generalization of the Erdős-Straus Conjecture
For the right-hand side to be strictly positive, it is necessary that $04x-n>$ 0, i.e., $x>\frac{n}{4}$, so we have $x\geq\left\lceil\frac{n}{4}\right\rceil$.
@@
for all $t$ in $\mathbb{N}^{*}$. If we replace $y+z$ and $yz$ with these expressions, the equation $yz(4x-n)=nx(y+z)$ will always hold.
Later, I will explain why I parameterized using $2t$ in this general case instead of using $t$ as I did for specific values of $n$. (Of course, we are allowed to parameterize using any multiple of $t$ in $\mathbb{N}^{*}$, as the equation will always hold.)
@@
$\Delta=4t^{2}(4x-n)^{2}-8ntx.$
Expanding this expression gives:
$\Delta=64t^{2}x^{2}-32t^{2}xn+4t^{2}n^{2}-8ntx.\quad(6)$
Now, let us determine for which values of $x$ we have $\Delta>0$. From (6), solving $\Delta>0$ is equivalent to solving:
$t(64tx^{2}-32txn+4tn^{2}-8nx)>0$
Since $t>0$, this simplifies to solving:
$064tx^{2}-32txn+4tn^{2}-8nx>$ 0 .
Solving this inequality gives:
$x\in\left]0,\frac{n(4t+1-\sqrt{8t+1})}{16t}\right[\cup\left]\frac{n(4t+1+\sqrt{8t+1})}{16t},+\infty\right[$
where $x$ belongs to $\mathbb{N}^{*}$.
To obtain integer solutions for this equation, it is necessary that $\Delta$ be a perfect square. I struggled to find an obvious perfect square for $\Delta$. It seems to me that the polynomial given in (6) is rich, but unfortunately, after some quick and unsuccessful attempts to find a classic perfect square of $\Delta$ without imposing any condition on $n$ (since we are in the process of generalization), the idea I had was to use the fact that the discriminant $\Delta$ holds for any $t\in\mathbb{N}^{*}$ and see what can be deduced to obtain a perfect square in order to bypass the difficulty. Let’s see how I proceed Let’s revisit the compact form of $\Delta$ in (5).
@@ 5 Compact Form of $\Delta$
This is valid for all $t\in\mathbb{N}^{*}$ because, as mentioned earlier, the parametrization is valid for all $t\in\mathbb{N}^{*}$. You also observed how this information allowed us to prove the conjecture (manually for some $n$) by providing specific values of $t$ that help us obtain a perfect square. We will use this information in this generalization.
@@
$x\in\left]0,\frac{n(4t+1-\sqrt{8t+1})}{16t}\right[\cup\left]\frac{n(4t+1+\sqrt{8t+1})}{16t},+\infty\right[$
@@
and since $\Delta>0$ for $x\in\left]0,\frac{n(4t+1-\sqrt{8t+1})}{16t}\right[\cup\left]\frac{n(4t+1+\sqrt{8t+1})}{16t},+\infty\right[$, for the x and t which respect this condition, it is clear that there exist $x$ and $t$ such that $t(4x-n)^{2}-2nx$ is in $\mathbb{N}^{*}$. We can then write:
@@
with $k\in\mathbb{N}^{*}$ Thus,
@@
which holds for all $t\in\mathbb{N}^{*}$. In particular, for $t=k$ (This is why we excluded the case $\Delta=0$, otherwise $k=\left(t(4x-n)^{2}-2nx\right)$ can be zero and we cannot set $t=k$ because from the start, during the parameterization, $t$ is non-zero). , so we have:
@@
We have thus obtained a perfect square. Let’s now see under what conditions this transformation holds.
@@ 6 Conditions for Obtaining a Perfect Square
This implies that:
$k=\frac{2nx}{(4x-n-1)(4x-n+1)},$
and $k\in\mathbb{N}^{*}$ if and only if $(4x-n-1)(4x-n+1)$ divides $2nx$. This means that both $(4x-n-1)$ divides $2nx$ and $(4x-n+1)$ divides $2nx$.
Under these conditions, $\Delta=4k^{2}=(2k)^{2}$, so:
@@
we have $y=k(4x-n-1)$ and $z=k(4x-n+1)$. Now, from earlier, we know that:
@@
Next, we find the conditions for $x$ such that both $y$ and $z$ are positive. Since $02nx>$ 0, we only need to ensure that the denominators are strictly positive.
@@
Therefore, $y\in\mathbb{N}^{*}$ if and only if $14x-n+$ 1 divides $2nx$, which implies:
@@
As we will see in the numerical checks, under these divisibility conditions, this formula will allow us to decompose the fraction $\frac{4}{n}$ into three unit fractions for those $n$ satisfying these conditions up to infinity (since we have shown that the values of $x$ for which $14x-n-$ 1 and $14x-n+$ 1 divide $2nx$ extend to infinity). This is good news because we always hope that for any given $n$, there is an infinite number of choices for $x$ that satisfy the divisibility condition, allowing us to prove the conjecture for that specific $n$.
@@
In summary, we have proven that for all $n\geq 2$, If there exists $x_{0}\in\mathbb{N}^{*}$ with $x_{0}\in\left[\left\lceil\frac{n+1}{4}\right\rceil+1,+\infty\right[$. such that $14x_{0}-n-$ 1 and $14x_{0}-n+$ 1 divide $2nx_{0}$, then there exist $y=\frac{2nx_{0}}{4x_{0}-n-1}$ and $z=\frac{2nx_{0}}{4x_{0}-n+1}$ in $\mathbb{N}^{*}$ such that:
@@
Below, I will provide numerical calculations that, depending on the range searched for $x_{0}$, cover a some $n$ Based on this formula verifying the conjecture, you will also be able to perform calculations using more powerful computational resources to identify very large $x_{0}$ and $n$. Of course, with numerical computations, it is impossible to reach +$\infty$+ ∞, but what assures us of the validity of the results is the analytical result we have just established. Every time we find such an $x_{0}$, the conjecture holds for the corresponding $n$.
@@
7 Second Formula: Condition of Existence of a Perfect Square
The first formula above is based on a divisibility condition, while the second one, which I will now introduce, relies on the existence of a perfect square. Let us reconsider the discriminant $\Delta$ of our quadratic equation:
$\Delta=4t^{2}(4x-n)^{2}-8ntx=4(t^{2}(4x-n)^{2}-2ntx).$
We have already seen that $\Delta\geq 0$ for $x\in\left]0,\frac{n\left(4t+1-\sqrt{8t+1}\right)}{16t}\right]\cup\left[\frac{n\left(4t+1+\sqrt{8t+1}\right)}{16t},+\infty\right[$ with $x\in\mathbb{N}^{*}$. Therefore, for $x$ that satisfies this condition, we define
$t^{2}(4x-n)^{2}-2ntx=q^{2},\quad q\in\mathbb{N}.$

### v2

# Partial Resolution of the Erdős–Straus, Sierpiński, and Generalized Erdős–Straus Conjectures Using New Analytical Formulas

@@ Abstract
This article proposes a unified analytical approach leading to a partial resolution of the Erdős–Straus, Sierpiński conjectures, and their generalization. We introduce an equivalent reformulation of these conjectures while constructing two new explicit analytical formulas. The first formula, which is a special case of the second, is based on a divisibility condition, whereas the second, more general formula, relies on the existence of a perfect square, which we conjecture to always hold. Under these conditions, the formulas verify the conjectures even for very large numerical values. Moreover, our method reduces the problem to the search for a suitable perfect square, thereby opening the way to a complete proof of these conjectures. In conclusion, we present open questions and conjectures to the mathematical community regarding the generalization of these formulas.
@@ 1 Introduction
The Erdős–Straus conjecture, formulated in 1948 by Paul Erdős and Ernst G. Strauss, states that for every natural number $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that
@@
Louis Mordell demonstrated that if $n$ is not congruent to $1,11^{2},13^{2},17^{2},19^{2},$ or $23^{2}$ modulo $840$, then the conjecture holds. Furthermore, numerical verifications have confirmed its validity up to $n=10^{17}$.
@@
The Polish mathematician Wacław Sierpiński also investigated this conjecture and posed a similar question: Is a similar decomposition possible if $4$ is replaced by $5$? He conjectured that for every natural number $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that
@@
Finally, following the formulation of the Erdős–Straus conjecture, several mathematicians explored the possibility of generalizing this Egyptian fraction representation. This led to the generalized Erdős–Straus conjecture, which may be more precisely formulated as follows: for every integer $a\geq 5$, there exists an integer $N_{0}\in\mathbb{N}$ with $N_{0}\geq 2$ (possibly depending on $a$, since it is easy to see that $N_{0}$ cannot be strictly smaller than $a/3$) such that for all $n\geq N_{0}$, there exist $x,y,z\in\mathbb{N}^{*}$ satisfying
@@
Moreover, this result is obtained using the same analytical formulas developed for the Erdős–Straus and Sierpiński conjectures.
@@ 2 The Erdős-Straus Conjecture
One of the methods that is known nowaday is that if $n$ is even integer , meaning $n=2k$ with $k\in\mathbb{N}^{*}$. In this case, it is well established that:
@@
In our new approach, I will demonstrate how this result can be naturally recovered when $n$ is even. Before introducing the general method that allowed me to establish two formulas verifying the conjecture under a certain condition, I will first illustrate my process by analytically testing the conjecture for some values of $n$.
Our method is based on the structural analysis of equation (1). It is a simple yet powerful and general approach, applicable to even numbers, odd numbers, and prime numbers alike.
@@ 2.1 Case $n=2$ (Even Number)
I choose not to simplify $\frac{4}{2}$ immediately in order to maintain the general structure of the method. However, **one can simplify this fraction directly**, and by following the same steps that I will present, the same result is always obtained.
@@
Since we are searching for solutions in $\mathbb{N}^{*}$, it is necessary that 4x24x-2 be strictly positive, that is, 4x2>04x-2>0, which gives $x\geq\lfloor\frac{1}{2}\rfloor+1=1$.
For the rest of the document, **the notation $\lfloor x\rfloor$ will always denote the integer part of xx** to avoid any confusion.
We then obtain the equation:
yz(4x2)=2x(y+z).yz(4x-2)=2x(y+z).
@@ 2.2 Parameterization Technique
Let $a,c\in\mathbb{R}$,and $b,d\in\mathbb{R^{*}}$ such that:
@@
This parametric representation holds for $t\in\mathbb{R}$ and allows rewriting $(dt)b=(bt)d$, which verifies the initial equation. This technique will be crucial for my results.
@@ 2.3 Application to Our Equation
This parameterization is valid for $t\in\mathbb{Q}{{}^{*}}_{+}$.
@@
For $y$ and $z$ to be integers, $\Delta$ must be a perfect square, and 4x24x-2 must have the same parity as $\sqrt{\Delta}$.
@@ 2.4 Analysis of the Method
One of the key advantages of our analytical approach is its applicability not only to even numbers but also to odd numbers and prime numbers. Unlike classical methods—which have relied primarily on brute-force techniques that merely test the existence of solutions without offering insight into the problem’s structure—our method derives solutions directly from the inherent logical structure of the underlying Diophantine equation. This provides a general, rigorous framework for addressing the Erdős–Straus, Sierpiński, and generalized Erdős–Straus conjectures. Moreover, the elegance and clarity of our approach make it a valuable pedagogical tool for teaching methods of solving Diophantine equations, as it not only yields the solutions but also deepens our understanding of the arithmetic properties involved
@@ 2.5 Case $n=3$ (odd and prime Number)
For $t=1$, we find a perfect square $11^{2}=121$, yielding $y=1$ and $z=12$.
@@ 2.7 Case $n=841$ (number not proven by Mordell)
NB. In most cases, we restrict the search for $t$ to $\mathbb{N}^{*}$. However, numerical simulations indicate that allowing $t$ to range over non-integer positive rationals $\mathbb{Q}_{+}^{*}\setminus\mathbb{N}$ becomes necessary for $a\geq 5$ when $n$ is small, namely in a neighborhood of the threshold $N_{0}$, since solutions remain sparse in this region. Beyond a certain threshold, we observe an abundance of integer values of $t$ for which the discriminant is a perfect square. Moreover, $t$ may also naturally take values in $\mathbb{Q}_{+}^{*}$ in the context of Formula (1) as soon as the divisibility condition is satisfied.
@@ 4 Generalization of the Erdős-Straus Conjecture
For the right-hand side to be strictly positive, it is necessary that 4xn>04x-n>0, i.e., $x>\frac{n}{4}$, so we have $x\geq\left\lceil\frac{n}{4}\right\rceil$.
@@
for $t\in\mathbb{Q}{{}^{*}}_{+}$. If we replace $y+z$ and $yz$ with these expressions, the equation yz(4xn)=nx(y+z)yz(4x-n)=nx(y+z) will always hold.
Later, I will explain why I parameterized using $2t$ in this general case instead of using $t$ as I did for specific values of $n$. (Of course, we are allowed to parameterize using any multiple of $t\in\mathbb{Q}{{}^{*}}_{+}$, as the equation will always hold.)
@@
To obtain integer solutions for this equation, it is necessary that $\Delta$ be a perfect square. I struggled to find an obvious perfect square for $\Delta$. It seems to me that the polynomial given in (5) is rich, but unfortunately, after some quick and unsuccessful attempts to find a classic perfect square of $\Delta$ without imposing any condition on $n$ (since we are in the process of generalization), the idea I had was to use the fact that the discriminant $\Delta$ holds for $t\in\mathbb{Q}_{+}^{*}$ arbitrary and see what can be deduced to obtain a perfect square in order to bypass the difficulty. Let’s see how I proceed Let’s revisit the compact form of $\Delta$ in (5):
@@ 5 Compact Form of $\Delta$
This is valid for $t\in\mathbb{Q}_{+}^{*}$ arbitrary because, as mentioned earlier, the parametrization is valid for $t\in\mathbb{Q}_{+}^{*}$ arbitrary. You also observed how this information allowed us to prove the conjecture (manually for some $n$) by providing specific values of $t$ that help us obtain a perfect square. We will use this information in this generalization.
@@
$t>\frac{2nx}{(4x-n)^{2}}.$
@@
and since $\Delta>0$ for
$t>\frac{2nx}{(4x-n)^{2}}.$
for the x and t which respect this condition, it is clear that there exist $x$ and $t$ such that $t(4x-n)^{2}-2nx$ is in $\mathbb{{Q}_{+}^{*}}$. We can then write:
@@
with $k\in\mathbb{Q}_{+}^{*}$ . Thus,
@@
which holds for $t\in\mathbb{Q}_{+}^{*}$ arbitrary. So, for $t=k$ (This is why we excluded the case $\Delta=0$, otherwise $k=\left(t(4x-n)^{2}-2nx\right)$ can be zero and we cannot set $t=k$ because from the start, during the parameterization, $t$ is non-zero). , so we have:
@@
We have thus obtained a perfect square (rational or integer perfect square). Let’s now see under what conditions this transformation holds.
@@ 6 Conditions for Obtaining a Perfect Square
This implies that
$k=\frac{2nx}{(4x-n-1)(4x-n+1)},\qquad k\in\mathbb{Q}_{+}^{*}.$
At this stage, no additional condition is required, since the above expression naturally defines a positive rational value of $k$ without any restriction.
Since, $\Delta=4k^{2}=(2k)^{2}$, so:
@@
we have y=k(4xn1)y=k(4x-n-1) and z=k(4xn+1)z=k(4x-n+1). Now, from earlier, we know that:
@@
A natural condition for $y$ and $z$ to be integers is that both 4xn14x-n-1 and 4xn+14x-n+1 divide $2nx$.
Next, we find the conditions for $x$ such that both $y$ and $z$ are positives. Since $2nx>0$, we only need to ensure that the denominators are strictly positive.
@@
Therefore, $y\in\mathbb{N}^{*}$ if and only if 4xn+14x-n+1 divides $2nx$, which implies:
@@
As we will see in the numerical checks, under these divisibility conditions, this formula will allow us to decompose the fraction $\frac{4}{n}$ into three unit fractions for those $n$ satisfying these conditions up to infinity (since we have shown that the values of $x$ for which 4xn14x-n-1 and 4xn+14x-n+1 divide $2nx$ extend to infinity). This is good news because we always hope that for any given $n$, there is an infinite number of choices for $x$ that satisfy the divisibility condition, allowing us to prove the conjecture for that specific $n$.
@@
In summary, we have proven that for all $n\geq 2$, If there exists $x_{0}\in\mathbb{N}^{*}$ with $x_{0}\in\left[\left\lceil\frac{n+1}{4}\right\rceil+1,+\infty\right[$. such that $4x_{0}-n-1$ and $4x_{0}-n+1$ divide $2nx_{0}$, then there exist $y=\frac{2nx_{0}}{4x_{0}-n-1}$ and $z=\frac{2nx_{0}}{4x_{0}-n+1}$ in $\mathbb{N}^{*}$ such that:
@@
we therefore have the following theorem:
Theorem 1.
Let $n\geq 2$ be an integer. Suppose there exists $x_{0}\in\mathbb{N}^{*}$ with
$x_{0}\in\left[\left\lceil\frac{n+1}{4}\right\rceil+1,+\infty\right)$
such that both integers $4x_{0}-n-1$ and $4x_{0}-n+1$ divide $2nx_{0}$. Then there exist $y,z\in\mathbb{N}^{*}$ given by
$y=\frac{2nx_{0}}{4x_{0}-n-1}\qquad\text{and}\qquad z=\frac{2nx_{0}}{4x_{0}-n+1},$
such that
$\frac{4}{n}=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}.$
Below, I will provide numerical calculations that, depending on the range searched for $x_{0}$, cover a some $n$ . Based on this formula verifying the conjecture, you will also be able to perform calculations using more powerful computational resources to identify very large $x_{0}$ and $n$. Of course, with numerical computations, it is impossible to reach $\infty$, but what assures us of the validity of the results is the analytical result we have just established. Every time we find such an $x_{0}$, the conjecture holds for the corresponding $n$.
@@
7 Second Formula: Condition of the Existence of a Perfect Square
While the first formula is derived from a divisibility condition, the second one is more fundamental in nature, as it provides a genuinely more general framework. Indeed, it is based on the existence of a perfect square and subsumes the first formula as a particular case. We now introduce this second formula.
Let us reconsider the discriminant $\Delta$ of our quadratic equation:
$\Delta=4t^{2}(4x-n)^{2}-8ntx=4\bigl(t^{2}(4x-n)^{2}-2ntx\bigr).$
We have already seen that $\Delta\geq 0$ for $t>\frac{2nx}{(4x-n)^{2}}$ with $t\in\mathbb{N}^{*}$. Therefore, for such values of $t$, we define
$t^{2}(4x-n)^{2}-2ntx=q^{2}$, with $q\in\mathbb{N}$ (as mentioned previously, here and in most cases we take $t$ and $q$ to be integers in order to automatically guarantee that, once the perfect square condition is satisfied, the resulting solutions are integral. Our parametrization allows $t$ to take arbitrary values without affecting the structure of the equation, and it is therefore up to us to choose suitable values of $t$ ensuring that the corresponding values of $y$ and $z$ are integers).